Prev Next

SW&IT Meeting Minutes 2022-03-07

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang


Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

LR20220221 - Bec, Matthieu 80.0
LR20220221 - Mechtley, Brandon 80.0
LR20220221 - Eychaner, Glenn 80.0
LR20220221 - Edmeades, David 24.0
LR20220221 - Grenz, Paul 40.0
LR20220221 - Hooper, Stephen 0.0
LR20220221 - Zhang, Yang 80.0
LR20220221 - Zhang, Xianyu 40.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports


  • Last week
    • OC, Management
    • ao-confcalib autmoated rev control deployed
    • LN tests
    • AO on sky efficiency task definitions (scope, resources, timeline, priority) meeting
    • AO SW requirements in the SOUL commissioning report and in the verification matrix meeting
    • SHARK-NIR: Risk & Hazard Assessment meeting
  • This week
    • ATT PM, chat with SE
    • set_z0 accumulator issue, pause-g0/offsets/skip
    • SHARK-NIR PSR documents review
    • safety training


  • Previous:
    • TCS and telescope support:
      • Worked with Paul on OSS/Proxy setup and tests (some problems remain).
      • Updated John Hill's new PMC configuration in TCS github repo.
      • Several support calls from Josh regarding OAC and AGW.
    • TMS:
      • Meeting and some slight work.
      • Remote shutter tests with mixed results (needs German team's inputs).
    • MODS/LBC:
      • Working on a new program for remote MODS data checking.
      • Meeting with Olga on MODS data checking/cleaning and DOFPIA.
      • IT reqest: Fedora VM for Olga with archive mount (more on this in the meeting).
    • No feedback yet from archive people.
  • Next:
    • Continue to support OSS/Proxy setup and tests.
    • Continue to work on MODS remote data checking program.
    • Follow-up meeting with Olga on various MODS/LBC issues.
    • Working with Olga on questions to OSU team related to IT#8567.
    • Resume some TMS software work and perhaps more TMS tests to come.


  • Previous
    • Quality of life stuff
      • Calendar invites to lists issue solved
      • Solidworks hiccup for MG
      • New laptop quote for AC
      • Vault 2020 licensing help for DR
      • Handholding IA2 people re: OSX sshd config for X forwarding/lbtplot
    • Sent strongly worded letter to Veeam contact, finally got reply. As suspected the new licensing scheme is much more expensive by a factor of 4.
      • At this point I have a number for 4 more (total 6) socket licenses: $2000.
      • I also asked for a number on a more full-featured socket license that allows direct cloud storage interaction, but I suspect it's going to be on the order of the above 4x cost.
    • Strange deactivation of DM's IT account.
      • Logging of such activity was not enabled, unfortunately, but it is now.
    • OK mentioned KVM certificate error; my initial suscpicion probably isn't the issue, but I'll investigate to see if there is a solution
    • Finally located RMA SSD, so we have a spare
  • Next
    • Veeam licensing direction
    • Test restore of Vault VM and its functionality from new backup setup.
      • If successful, I can remove a significant chunk of data from Backblaze
    • Is it time to talk to archive about the newdata volume swap?


  • Last week:
    • IT 8579 (Passive Flexure Correction): After more investigation, implemented a fix (LUCI 2.14.7) suggested by D.Thompson; see the notes for details.
    • Finally got a chance to look at some performance tests I did weeks ago, and found some issues with the GUI overloading the Event Dispatch thread, Executor thread, and Aladin command processor. Deployed a new RTD (2.14.96) which should further improve the performance issues.
  • This week
    • Keep testing the new RTD and fix minor bugs as they come up.
    • Restart work on the observer GUI.
    • Clean up the scripts in lcsp/bin on luci.luci and in stow on observer machines. They have drifted significantly from what is checked into GitHub.
    • I will be out Tuesday morning for a medical appointment.


  • Last week:
    • rescheduling set_z0 bank tests, pending more discussion with Guido/Xianyu or some consensus?
    • interrupted offset investigation with Juan Carlos
    • LN testing
    • AOIce fix--default ports
    • aoarbitrator--looking into which commands observe interrupt, which enable/disable interrupt and cancel buttons
  • This week:
    • resume/gain ramp order--error handling if either fsm_set_gain fails.
    • MsgD issues:
      • AOS disconnect from this weekend? (Doug)
      • AOS dropping notification of wfsarb.L.MSG, doesn't always show messages that disturbances have loaded.
    • interrupted offsets (saved modes)
    • autogain interrupted by loopfault--set gains to 1


  • Last week
    • Finished NOMIC code updates for FITS header.
    • Help with MAPS camera code.
    • Help with SHARK-NIR
    • Testing proxy for white light source
    • Argos software won't connect properly with proxy connected; more tests are needed
  • This week
    • Testing NOMIC
    • Testing proxy code to see why connections are failing
    • Examining GCS


  • Moving data from old to new Synology
  • DIMM
  • Mountain network stuff
  • VMware machine
  • IT meeting

  • VMware installation mountain
  • Upgrade 507 computer


  • Last week
    • Mixed mode control SW test on the mountain
  • This week
    • Mixed mode control test on the mountain
Topic revision: r11 - 07 Mar 2022, XianyuZhang
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback